OpenStudy (anonymous):

Evaluate and write your answer in scientific notation:(7 x 10-2)(1.5 x 10-3)

OpenStudy (anonymous):

Well. 7*1.5 = 10.5 And 10^-2 * 10^-3 = 10^(-2+-3) = 10^-5 so its 10.5*10^-5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem